# Ctrl-Room: Controllable Text-to-3D Room Meshes Generation with Layout Constraints

### v3

@@ 8 Scene Code Denoising Network
In the Layout Generation Stage, we use a holistic scene code to parameterize the indoor scene and design a diffusion model to learn its distribution. Specifically, given a 3D scene $\mathcal{S}$ with $N$ objects, we represent the scene layout as a holistic scene code $\mathbf{x_{0}}=\{\mathbf{o_{i}}\}^{N}_{i=1}$. We encode each object $o_{i}$ as a node with various attributes, i.e., center location $l_{i}\in\mathbb{R}^{3}$, size $s_{i}\in\mathbb{R}^{3}$, orientation $r_{i}\in\mathbb{R}$, class label $c_{i}\in\mathbb{R}^{C}$. Each node is characterized by the concatenation of these attributes as $\mathbf{o}_{i}=[c_{i},l_{i},s_{i},r_{i}]$. As shown in Fig. 10, our scene code denoising network of the layout diffusion model is built upon IDDPM [21]. The whole architecture of the layout diffusion model is similar to IDDPM, while we replace the upsample and downsample blocks with 1D-convolution network in the U-Net, and insert attention blocks after each residual block to capture both the global context among objects and the semantic context from the input text prompt. The input encoding head processes different encoding of the node attributes, e.g., semantic class labels, box centroid, and box orientation. After adding noise, the input encoding is fed into the U-Net to obtain a denoised scene code. During the forward phase, as in IDDPM, we iteratively perform the denoising process and generate a scene code from a partial scene textual description.
@@ 9 Mask-Guided Editing
To achieve consistent and seamless 3D scene editing, it should achieve two goals, i.e. altering the content according to the user’s input, and maintaining appearance consistency for scene objects. We propose a mask-guided image editing as illustrated Fig.4 in the main paper, where a chair’s position is moved. In the following, we will explain our method with this example. We denote the semantic panorama from the edited scene as $P_{\rm edited}$, then we derive the guidance masks based on its difference from the original one $P_{\rm ori}$. The source mask $\mathbf{m}_{\rm src}$ shows the position of the original chair, and the target mask $\mathbf{m}_{\rm tar}$ indicates the location of the moved chair, and the inpainting mask $\mathbf{m}_{\rm inpaint}=\{m|m\in\mathbf{m}_{\rm src}\;{\rm and}\;m\notin\mathbf{m}_{\rm tar}\}$ is the unoccluded region. Given these guidance masks, our method includes two steps: the inpainting step and the optimization step. We first fill in the inpaint area by feeding the inpaint mask $\mathbf{m}_{\rm inpaint}$ and edited semantic panorama $P_{\rm edited}$ to the inpainting step. Then, in our optimization step, we optimize the DIFT [38] feature to maintain the visual consistency of relocated objects.
@@
$\mathbf{x_{t}^{\rm ori}}$ $\sim$ $\mathcal{N}(\sqrt{\bar{\alpha_{t}}}\mathbf{x_{0}^{\rm ori}},(1-\bar{\alpha_{t}}\mathbf{I})),$ (7)
$\mathbf{x_{t}^{{\rm new}}}$ $\sim$ $\mathcal{N}(\mu_{\theta}(x_{t},t,y,P_{\rm edited}),\Sigma_{\theta}(x_{t},t,y,P_{\rm edited})),$ (8)
$\hat{\mathbf{x}}_{t-1}^{\rm new}$ $=$ $\mathbf{m}_{\rm inpaint}\odot\mathbf{x_{t}^{{\rm new}}}+(1-\mathbf{m}_{\rm inpaint})\odot\mathbf{x_{t}^{\rm ori}},$ (9)
where $\mathbf{x^{{\rm ori}}_{t}}$ is obtained through propagating $\mathbf{x^{{\rm ori}}_{0}}$ in diffusion process, and $\mathbf{x_{t}^{{\rm new}}}$ is sampled from the fine-tuned ControlNet model, which takes the edited semantic layout panorama $P_{\rm edited}$ and text prompt $y$ as input. As the propagated $\mathbf{x^{{\rm ori}}_{t}}$ is unaware of the new content $\mathbf{x_{t}^{{\rm new}}}$, this may result in distracting boundaries of the inpainted area. To better blend the new content $\mathbf{x}^{\rm new}_{t}$ and its surrounding background $\mathbf{x^{{\rm ori}}_{t}}$ in the inpainted area, we update the computation of $\hat{\mathbf{x}}_{t-1}^{\rm new}$ to,
$\hat{\mathbf{x}}_{t-1}^{\rm new}=\mathbf{m}_{\rm inpaint}\odot\mathbf{x_{t}^{{\rm new}}}+(1-\mathbf{m}_{\rm inpaint})\odot(\mathbf{x_{t}^{\rm ori}}\cdot\lambda_{\rm ori}+\mathbf{x_{t+1}^{\rm new}}\cdot\lambda_{\rm new}),$ (10)
@@
$\mathcal{L_{\rm obj}}=-cos({\rm sg}(F^{ori}_{t}\odot\mathbf{m}_{\rm src}),F^{edit}_{t}\odot\mathbf{m}_{\rm tar}).$ (11)
@@
$\hat{\mathbf{x}}^{k+1}_{t}=\hat{\mathbf{x}}^{k}_{t}-{\eta}\cdot\frac{\partial\mathcal{L_{\rm obj}}}{\partial\hat{\mathbf{x}}^{k}_{t}}.$ (12)
@@ 10 Dataset
  • data filtering. We start by filtering out those problematic scenes such as rooms with wall number less than $4$ or larger than $20$. We also remove those scenes with too few or too many objects. The number of walls of valid bedrooms is between $4$ and $10$, and that of objects is between $3$ and $13$. As for living rooms, the minimum and maximum numbers of walls are set to $4$ and $20$, and that of objects are set to $3$ and $24$ respectively. The number of walls for valid kitchens, studies, and bathrooms is the same as for bedrooms, while the objects number is between $3$ and $24$. Thus, the number of scene nodes is $N=23$ in bedrooms, $N=44$ in living rooms, and $N=34$ in kitchens, studies, and bathrooms. After filtering, we get 4,961 bedrooms, 3,039 living rooms, 1,848 kitchens, 638 studies, and 1,356 bathrooms.
Figure 12: Example of object bounding box annotation.
@@ 11 Implementation details
Table 3: The computational cost comparison on A6000. Since Text2Room does not offer a standardized neural network, we cannot measure its parameters.
Method Inference Time/s GPU Memory Params
Text2Light 81.56 5.46G 630.66M
MVDiffusion 208.5 8.74G 1352.54M
Text2Room $>9,000$ $>16$G -
Ours 61.1 1.95G + 10.41G 63.51M + 1220.62M
@@ 11.1 Baseline Implementations
  • MVDiffusion [39]: To get a high-resolution photo realistic panorama, MVDiffusion employs $8$ branches of SD [25] model and correspondence-aware attention mechanism to generate multi-view images simultaneously. We first fine-tune the pre-trained model of MVDiffusion on Structured3D for 10 epochs(about 3 days). Since each generated subview image of MVDiffusion is at $512\times 512$ resolution, the final panorama is pretty large. We resize the generated panorama of MVDiffusion from $4096\times 2048$ to $1024\times 512$. Then the 8 subview perspective images are extracted from the post-processed panorama using the same camera settings (FOV=$90^{\circ}$,rotation=$45^{\circ}$). The same operation is adopted on our generated panoramic images. Finally, we combine the panorama from MVDiffusion with our panoramic reconstruction module to create a 3D mesh.
  • Text2Light [6]: Text2Light creates HDR panoramic images from text using a multi-stage auto-regressive generative model. We choose Text2Light as one of the baseline for our panorama generation and 3D room mesh generation. We first generate RGB panoramas from the input text using Text2Light, then lift it into 3D mesh using the same panoramic reconstruction module as our method. When evaluate the panoramic image quality, we adopt the same processing as MVDiffusion to get multi-view perspective images of Text2Light.
@@
We also compare the model complexity between ours and the baselines, Table 3 shows our method achieves significantly faster runtime for generating a 3D room compared to existing methods. The difference in GPU memory footprint and parameter quantity is not significant. Despite having two diffusion models, our layout generation stage only needs to produce high-level room layouts and furniture arrangements, resulting in low computational costs and model complexity.
@@ 12 Panorama Generation Comparison
Fig. 13 presents additional results for panorama generation. Given a simple partial-scene text prompt, our approach obtains better RGB panorama than that of Text2Light [6] and MVDiffusion [39], which demonstrates the effectiveness of our well-designed framework. While Text2Light suffers from the inconsistent loop and unexpected content of the generated panorama, MVDiffusion fails to recover a reasonable room layout from the text prompt.
@@ 13 Additional Qualitative Results
Figure 16: Additional qualitative comparison with previous works.
@@ 15 Free style prompts
Describe what is displayed in the panoramic image succinctly in 3 or 4 sentences encoded in ASCII. Do not use lengthy or compound sentences. Do not mention that it is an image or a panoramic image. Do not describe the background, lighting, color palette or count the number of objects. Do not describe size like “small”, “large”, etc. Describe the relative positions of each objects in the scene using only these relationships: “on”, “above”, “surrounding”, “inside”, “left touching”, “right of”, “front touching”, “in front of”, “right touching”, “left of”, “behind touching”, “behind”, “next to”, “left of”, “right of”. Optionally, describe the object attributes (color, texture etc). In the description only use these objects: table, night stand, picture, door, cabinet, curtain, bathtub, bed, sink, fridge, shelves, window, lamp, chair, pillow, dresser, bookshelf, sofa, counter, desk, mirror, television, wall

### v4

@@ 7 Scene Code Denoising Network
In the Layout Generation Stage, we use a holistic scene code to parameterize the indoor scene and design a diffusion model to learn its distribution. Specifically, given a 3D scene $\mathcal{S}$ with $N$ objects, we represent the scene layout as a holistic scene code $\mathbf{x_{0}}=\{\mathbf{o_{i}}\}^{N}_{i=1}$. We encode each object $o_{i}$ as a node with various attributes, i.e., center location $l_{i}\in\mathbb{R}^{3}$, size $s_{i}\in\mathbb{R}^{3}$, orientation $r_{i}\in\mathbb{R}$, class label $c_{i}\in\mathbb{R}^{C}$. Each node is characterized by the concatenation of these attributes as $\mathbf{o}_{i}=[c_{i},l_{i},s_{i},r_{i}]$. As shown in Fig. 9, our scene code denoising network of the layout diffusion model is built upon IDDPM [20]. The whole architecture of the layout diffusion model is similar to IDDPM, while we replace the upsample and downsample blocks with 1D-convolution network in the U-Net, and insert attention blocks after each residual block to capture both the global context among objects and the semantic context from the input text prompt. The input encoding head processes different encoding of the node attributes, e.g., semantic class labels, box centroid, and box orientation. After adding noise, the input encoding is fed into the U-Net to obtain a denoised scene code. The training objectives includes the denoising objective $\mathcal{L_{\rm denoise}}$ and a regularization term $\mathcal{L}_{\rm physical}$ to penalize the penetration among objects and walls as follows,
$\mathcal{L}$ $=$ $\mathcal{L_{\rm denoise}}+\mathcal{L}_{\rm physical},$ (5)
$\begin{split}\mathcal{L}_{\rm\mathbf{w-o}}&=\sum_{i=1}^{K_{\rm wall}}\sum_{j=1}^{K_{\rm object}}\sum_{p=1}^{8}Relu[-(a_{i}x_{jp}+b_{i}y_{jp}+c_{i}z_{jp}+d_{i})]\\ &*\mathds{1}(\prod_{\mathbf{w}_{i}}(x_{jp},y_{jp},z_{jp})\,{\rm in}\,\mathbf{w}_{i}).\end{split}$ (8)
Here, $(a_{i},b_{i},c_{i})$ is the normal vector of wall $\mathbf{w}_{i}$ that points towards the room center. $\prod_{\mathbf{w}_{i}}$ is the operator projecting a point onto the plane defined by $\mathbf{w}_{i}$. The plane equation of i-th wall is $a_{i}x+b_{i}y+c_{i}z+d=0$ and $\mathds{1}(\prod_{\mathbf{w}_{i}}(x_{jp},y_{jp},z_{jp})\,{\rm in}\,\mathbf{w}_{i})$ indicates whether the projection of bounding box vertices $(x_{jp},y_{jp},z_{jp})$ of j-th object is inside $\mathbf{w}_{i}$. We skip some objects such as windows and doors since they can intersect with walls. We adopt the 3D IoU loss $\mathcal{L}_{\rm o-o}$ in DiffuScene [34] as follows,
During the forward phase, as in IDDPM, we iteratively perform the denoising process and generate a scene code from a partial scene textual description.
@@
Figure 11: Our Layout-guided depth estimation. To align the estimated depth map with the 3D scene layout, we render depth map and instance map from the 3D scene layout at current view. Then we take the background depth (wall, ceiling, floor) as reference to align the depth prediction by optimizing the pretrained MDE. Avoiding degrade the MDE’s performance at object surface, we ensure the normal consistency of each object during the optimization process.
8 Layout-Guided Panoramic NeRF
Since the generated panorama is a partial observation of the room subject to occlusions, lifting the single view into a 3D room becomes a complex problem. Here we adopt the Warp and Inpainting scheme to complete the 3D room progressively. After generating the panorama $I^{0}$, we recover its depth map $D^{0}$ using the state-of-the-art monocular depth estimator (MDE)  [45]. However, problems such as scale ambiguity and large occlusions may lead to incomplete 3D room reconstruction. Additionally, ensuring consistency in inpainted panoramas at new viewpoints poses another challenge. Fortunately, the scene layout generated in the first stage offers crucial geometric and semantic guidance, which can help correct biased depth predictions and guide the inpainting process to generate novel view panoramas. In this paper, we employ PeRF [38] as the 3D room model and progressively generate novel viewpoint panoramas with layout guidance to reconstruct the PeRF model.
Figure 12: The Layout-guided PeRF takes the input panorama, aligned depth map and normal map as initialization. Then a progressive inpainting module is introduced to generate consistent panoramic images at the sampled novel views. The progressive inpainting module consists of the layout-guided panorama inpainting and the layout-guided depth estimation module. The final RGB-D panoramic pairs are included as training views to finetune PeRF [38].
Layout-Guided Depth Estimation.
To align the estimated panoramic depth map with the scene layout, a naive approach would be to directly compute scale and bias coefficients for the initial depth map $D^{0}$. However, as the scene layout consists of object bounding boxes and can not provide pixel-level perfect depth supervision, this method may lead to degraded depth predictions. To address this problem, we propose the panoramic geometry alignment module as depicted in Fig. 11. We use the instance labels of furniture items to exclude the rendered depths within the furniture areas, retaining only the background depth map (e.g., wall, ceiling, floor) denoted as $D^{0}_{l}$. We incorporate a consistency loss $\mathcal{L}_{\rm align}$ to optimize the pre-trained monocular depth estimator (MDE). This consistency loss is formulated as follows,
$\mathcal{L}_{\rm align}$ $=$ $\mathcal{L_{\rm d}}*w_{d}+\mathcal{L}_{\rm n}*w_{n},$ (10)
$\mathcal{L_{\rm d}}$ $=$ $L_{1}(D^{0}_{i},D^{0}_{l})*(\sim M^{0}),$ (11)
$\mathcal{L}_{\rm n}$ $=$ $\left|N^{0}-N^{0}_{i}\right|*M^{0}.$ (12)
where $\mathcal{L}_{\rm d}$ represents the smooth L1 loss of depth, $\mathcal{L}_{\rm n}$ denotes the absolute loss of normal, and $w_{d}$, $w_{n}$ are weighting coefficients. $D^{0}_{i}$ stands for the predicted depth of the MDE at i-th iteration. $M^{0}$ is the instance map denoting the foreground, such that we only correct the predicted depth in the background region while preserving normals in the furniture regions. After alignment, we obtain the optimal depth map $D^{*}$ and normal map $N^{*}$.
Layout-Guided Novel View Generation. PeRF [38] trains a panoramic neural radiance field using a single panorama. To render novel view panoramas, it employs the Stable Diffusion model [24] to inpaint 60 perspective images and stitch them into a panorama. Although it produces consistent renderings at nearby viewpoints, it struggles to synthesize viewpoints that are far apart and involve large unoccluded regions. To address this limitation, we rely on the scene layout to guide the panorama inpainting to maintain cross-view consistency.
Figure 13: After the layout-guided panorama inpainting, Our generated panoramas at novel viewpoints adhere to the semantic layout and seamlessly integrate with the visible regions, while PeRF [38] fails to synthesize plausible content at those large-size occlusion areas.
As illustrated in Fig. 12, we initialize the scene NeRF with the panorama $I^{0}$, the aligned depth map $D^{*}$ and normal map $N^{*}$. We sample new viewpoints in the green area of the occupancy grid that do not conflict with the initial furniture arrangement. At the i-th novel view, we render semantic map $S^{i}_{l}$, depth map $D^{i}_{l}$ and instance map $M^{i}_{l}$ from the scene layout, these are then combined with the panoramic rendering $I^{i}_{r}$ and inpainting mask $\mathbf{m}_{\rm inpaint}$ obtained from the NeRF and fed to the layout-guided panorama inpainting module to generate the novel view panorama. Using our fine-tuned ControlNet, it achieves training-free panoramic inpainting, which replaces pixels outside the inpainting mask $\mathbf{m}_{\rm inpaint}$ with $I^{i}_{r}$ and fill $\mathbf{m}_{\rm inpaint}$ based on the semantic map $S^{i}_{l}$. As demonstrated in Fig. 13, our resulting RGB panorama adheres to the semantic layout and seamlessly integrates with the visible regions, while that of PeRF [38] fails to generate reasonable content in the large occlusion areas. Subsequently, after generating the novel view panorama, we apply the layout-guided depth estimation and include it as training views for PeRF following their framework [38].
@@ 9 Mask-Guided Editing
To achieve consistent and seamless 3D scene editing, it should achieve two goals, i.e. altering the content according to the user’s input, and maintaining appearance consistency for scene objects. We propose a mask-guided image editing as illustrated in Fig. 14, where a chair’s position is moved. In the following, we will explain our method with this example. We denote the semantic panorama from the edited scene as $S_{\rm edited}$, then we derive the guidance masks based on its difference from the original one $S_{\rm ori}$. The source mask $\mathbf{m}_{\rm src}$ shows the position of the original chair, and the target mask $\mathbf{m}_{\rm tar}$ indicates the location of the moved chair, and the inpainting mask $\mathbf{m}_{\rm inpaint}=\{m|m\in\mathbf{m}_{\rm src}\;{\rm and}\;m\notin\mathbf{m}_{\rm tar}\}$ is the unoccluded region. Given these guidance masks, our method includes two steps: the inpainting step and the optimization step. We first fill in the inpaint area by feeding the inpaint mask $\mathbf{m}_{\rm inpaint}$ and edited semantic panorama $S_{\rm edited}$ to the inpainting step. Then, in our optimization step, we optimize the DIFT [35] feature to maintain the visual consistency of relocated objects.
@@
$\mathbf{x_{t}^{\rm ori}}$ $\sim$ $\mathcal{N}(\sqrt{\bar{\alpha_{t}}}\mathbf{x_{0}^{\rm ori}}(1-\bar{\alpha_{t}}\mathbf{I})),$ (13)
$\mathbf{x_{t}^{{\rm new}}}$ $\sim$ $\mathcal{N}(\mu_{\theta}(x_{t},t,y,S_{\rm edited}),\Sigma_{\theta}(x_{t},t,y,S_{\rm edited}))$ (14)
$\hat{\mathbf{x}}_{t-1}^{\rm new}$ $=$ $\mathbf{m}_{\rm inpaint}\odot\mathbf{x_{t}^{{\rm new}}}+(1-\mathbf{m}_{\rm inpaint})\odot\mathbf{x_{t}^{\rm ori}}$ (15)
where $\mathbf{x^{{\rm ori}}_{t}}$ is obtained through propagating $\mathbf{x^{{\rm ori}}_{0}}$ in diffusion process, and $\mathbf{x_{t}^{{\rm new}}}$ is sampled from the fine-tuned ControlNet model, which takes the edited semantic layout panorama $S_{\rm edited}$ and text prompt $y$ as input. As the propagated $\mathbf{x^{{\rm ori}}_{t}}$ is unaware of the new content $\mathbf{x_{t}^{{\rm new}}}$, this may result in distracting boundaries of the inpainted area. To better blend the new content $\mathbf{x}^{\rm new}_{t}$ and its surrounding background $\mathbf{x^{{\rm ori}}_{t}}$ in the inpainted area, we update the computation of $\hat{\mathbf{x}}_{t-1}^{\rm new}$ to,
$\begin{split}\hat{\mathbf{x}}_{t-1}^{\rm new}&=\mathbf{m}_{\rm inpaint}\odot\mathbf{x_{t}^{{\rm new}}}\\ &+(1-\mathbf{m}_{\rm inpaint})\odot(\mathbf{x_{t}^{\rm ori}}\cdot\lambda_{\rm ori}+\mathbf{x_{t+1}^{\rm new}}\cdot\lambda_{\rm new})\end{split}$ (16)
@@
$\mathcal{L_{\rm obj}}=-cos({\rm sg}(F^{ori}_{t}\odot\mathbf{m}_{\rm src}),F^{edit}_{t}\odot\mathbf{m}_{\rm tar})$ (17)
@@
$\hat{\mathbf{x}}^{k+1}_{t}=\hat{\mathbf{x}}^{k}_{t}-{\eta}\cdot\frac{\partial\mathcal{L_{\rm obj}}}{\partial\hat{\mathbf{x}}^{k}_{t}}$ (18)
@@ 10 Dataset
  • data filtering. We start by filtering out those problematic scenes such as rooms with wall number less than $4$ or larger than $20$. We also remove those scenes with too few or too many objects. The number of walls of valid bedrooms is between $4$ and $10$, and that of objects is between $3$ and $13$. As for living rooms, the minimum and maximum numbers of walls are set to $4$ and $20$, and that of objects are set to $3$ and $24$ respectively. The number of walls for valid kitchens, studies, and bathrooms is the same as for bedrooms, while the objects number is between $3$ and $24$. Thus, the number of scene nodes is $N=23$ in bedrooms, $N=44$ in living rooms, and $N=34$ in kitchens, studies, and bathrooms. After filtering, we get 4,961 bedrooms, 3,039 living rooms, 1,848 kitchens, 638 studies, and 1,500 bathrooms.
@@ 11 Implementation details
  • As for the layout-guided panoramic NeRF module, we set $w_{d}=0.6$ and $w_{n}=0.4$ for depth alignment loss. During the NeRF fitting process, we randomly select 8 viewpoints for living room scenarios and 4 viewpoints for other room types. The NeRF training settings are the same as PeRF [38].
@@ 11.1 Baseline Implementations
  • MVDiffusion [36]: To get a high-resolution photo realistic panorama, MVDiffusion employs $8$ branches of SD [24] model and correspondence-aware attention mechanism to generate multi-view images simultaneously. We first fine-tune the pre-trained model of MVDiffusion on Structured3D for 10 epochs(about 3 days). Since each generated subview image of MVDiffusion is at $512\times 512$ resolution, the final panorama is pretty large. We resize the generated panorama of MVDiffusion from $4096\times 2048$ to $1024\times 512$. Then the 8 subview perspective images are extracted from the post-processed panorama using the same camera settings (FOV=$90^{\circ}$,rotation=$45^{\circ}$). The same operation is adopted on our generated panoramic images. Finally, we combine the panorama from MVDiffusion with the depth estimation [45] and Poisson reconstruction [14] module to create a 3D mesh.
  • Text2Light [6]: Text2Light creates HDR panoramic images from text using a multi-stage auto-regressive generative model. We choose Text2Light as one of the baseline for our panorama generation and 3D room mesh generation. We first generate RGB panoramas from the input text using Text2Light, then lift it into 3D mesh using the same panoramic reconstruction module as MVDiffusion. When evaluating the panoramic image quality, we adopt the same processing as MVDiffusion to get multi-view perspective images of Text2Light.
@@
To ensure a fair comparison, we render 60 perspective images at the origin using the final textured meshes of all methods. The camera field of view is set to $140^{\circ}$ to capture scene layout information for evaluating the CS and IS scores. Additionally, we render corresponding geometric images in Fig. 20 to showcase the geometry quality.
@@ 12 Panorama Generation Comparison
In Fig. 16, we study the performance of our panorama generation module with and without loop-consistent sampling mechanism, the ablation indicates the loop-consistent sampling helps the generated panorama obtain better texture consistency. Fig. 17 presents additional results for panorama generation. Given a simple partial-scene text prompt, our approach obtains better RGB panorama than that of Text2Light [6] and MVDiffusion [36], which demonstrates the effectiveness of our well-designed framework. While Text2Light suffers from the inconsistent loop and unexpected content of the generated panorama, MVDiffusion fails to recover a reasonable room layout from the text prompt.
@@ 13 Additional Qualitative Results
Figure 20: Additional qualitative comparison with previous works.The first row shows a textured 3D room model, and the second row shows perspective colored renderings and geometric renderings from the room model.
@@ 15 Free style prompts
Describe what is displayed in the panoramic image succinctly in 3 or 4 sentences encoded in ASCII.
Do not use lengthy or compound sentences. Do not mention that it is an image or a panoramic image.
Do not describe the background, lighting, color palette or count the number of objects.
Do not describe size like “small”, “large”, etc.
Describe the relative positions of each objects in the scene using only these relationships: “on”, “above”, “surrounding”, “inside”, “left touching”, “right of”, “front touching”, ‘in front of”, “right touching”, “left of”, “behind touching”, “behind”, “next to”, “left of”, “right of”. Optionally, describe the object attributes (color, texture etc).
In the description only use these objects: table, night stand, picture, door, cabinet, curtain, bathtub, bed, sink, fridge, shelves, window, lamp, chair, pillow, dresser, bookshelf, sofa, counter, desk, mirror, television, wall